3. The position of a particle moving along the x-axis is given by x = 7t^2 + 2t^3, where x is in meters and t is in seconds. Determine the following, when t = 1.4 s:

(a) The position (2) (b) The velocity (2) (c) The acceleration (2)


1
Expert's answer
2022-03-24T16:18:46-0400

a)

"x=7t^2+2t^3,"

"x(1.4)=19.2~m,"

b)

"v=\dot x=14t+6t^2,"

"v(1.4)=31.4~\frac ms,"

c)

"a=\dot v=14+12t,"

"a(1.4)=30.8~\frac m{s^2}."